OpenStudy (anonymous):

There are eight seniors and six juniors in the Math Club at Jefferson High School. In how many ways can a math team consisting of four seniors and three juniors be selected from members of the Math Club? what is the answer i'm doing a test right now

OpenStudy (kropot72):

The required number of ways is found as follows: 8C4 * 6C3=?

OpenStudy (anonymous):

howdo i do that?

OpenStudy (anonymous):

i got1400

OpenStudy (kropot72):

\[\frac{8!}{4!4!}\times \frac{6!}{3!3!}=70\times 20\]